### v1

# Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model

Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model
We present results from the latest version of the GAlaxy Evolution and Assembly (gaea ) theoretical model of galaxy formation coupled with merger trees extracted from the P-Millennium Simulation (PMS ). With respect to the Millennium Simulation, typically adopted in our previous work, the PMS provides a better mass resolution ( 10), a larger volume (8003 Mpc3) and assumes cosmological parameters consistent with latest results from the Planck mission. The model includes, at the same time, a treatment for the partition of cold gas into atomic and molecular (H2) components; a better treatment for environmental processes acting on satellite galaxies; an updated modelling of cold gas accretion on Super-Massive Black Hole, leading to the Active Galactic Nuclei (AGN) phenomenon and relative feedback on the host galaxy. We compare gaea predictions based on the PMS , with model realizations based on other simulations in the Millennium Suite at different resolution, showing that the new model provides a remarkable consistency in most statistical properties of galaxy populations. We interpret this as due to the interplay between AGN feedback and H2-based SFR (both acting as regulators of the cold gas content in model galaxies), as model version considering only one of the two mechanisms do not show the same level of consistency in the predictions. We then compare model predictions with available data for the galaxy 2-point correlation function (2pCF) in the redshift range 0¡z3. We show that gaea runs are able to correctly recover the main dependencies of the 2pCF as a function of stellar mass (M⋆), star formation activity, HI-content and redshift for M⋆ ¡ 1011 M⊙ galaxies. These results suggest that our model correctly captures both the distribution of galaxy populations in the Large Scale Structure and the interplay between the main physical processes regulating their baryonic content, both for central and satellite galaxies. At larger stellar masses gaea underpredicts the 2pCF amplitude, suggesting that model massive galaxies live in less massive dark matter haloes. The model predicts a rather small redshift evolution of the clustering amplitude up to z3, consistent with available observational evidence.
Key Words.:
galaxies: formation – galaxies: evolution – galaxies: star formation – galaxies: statistics – galaxies: stellar content1 Introduction
The evolution of the large-scale spatial distribution of galaxies along cosmic epochs, i.e. the so-called Large Scale Structure (LSS) of the Universe, provides one of the most powerful constraints in observational cosmology. Galaxy surveys (see, e.g., Sánchez et al., 2006) have also shown the relevance of galaxy correlation functions, and in particular the two-point correlation function (2pCF), as a useful statistical estimator to characterize the LSS. The redshift evolution of galaxy 2pCF seems to be weaker/slower that the expected evolution of the clustering for the underlying mass distribution in concordance cosmologies (Jenkins et al., 1998): this mismatch is usually interpreted as the effect of physical mechanisms acting on the baryonic component. This implies that the redshift evolution of the 2pCF can also provide important constraints for baryon physics and galaxy formation models. Indeed, successful theoretical models of galaxy evolution on cosmological volumes usually aim to reproduce the redshift evolution of key statistical estimators, which are revealing of the correct build up of the stellar mass (M) as a function of redshift, like the galaxy stellar mass function (GSMFs) or luminosity functions (LFs). Nonetheless, it is also of fundamental importance that these models are able to correctly describe how the different galaxy types populated host dark matter (DM) haloes (and sub-haloes) in different density environments (see, e.g., Cole & Kaiser, 1989).
According to the classical interpretation of the 2pCF shape, it can be seen as a combination of (see, e.g., Cooray & Sheth, 2002): a 2-halo term - corresponding to pairs of galaxies belonging to independent halos - and a 1-halo term - coming from pairs of galaxies living in the same halo. The 2-halo term is dominant at large scales (separations larger than 3 Mpc) and comes mainly from pairs of central galaxies. The 1-halo term is responsible for the 2pCF shape at 2 Mpc and it is statistically dominated by galaxy pairs belonging to the same halo (that is at least one of the two galaxies is a satellite). Overall, the largest contribution to the global 2pCF comes from galaxies around the knee of the GSMF, i.e. with stellar masses similar to the Milky-Way (Li & White, 2009). The small scale clustering is particularly sensitive to the number and distribution of satellite galaxies inside massive DMHs. Therefore, it provides important constraints to theoretical models of galaxy formation, as it critically depends on the treatment adopted for satellite galaxies and in particular their survival times. Therefore, it can be significantly affected by numerical resolution effects (Guo & White, 2014).
Observationally, only the projected 2pCF - - is directly accessible, while the tridimensional 2pCF can be recovered via de-projection. Early results by Davis & Geller (1976) already showed a clear dependence of clustering amplitude with galaxy type, by showing that early type galaxies are more clustered than late-type galaxies. The advent of large area redshift surveys like the Sloan Digital Sky Survey (SDSS) provided the statistics to explore the dependence of amplitude and shape on galaxy properties such as luminosity (see, e.g., Zehavi et al., 2002) or stellar mass (Li et al., 2006), showing that brighter/more massive galaxies are more clustered than their fainter/less massive counterparts. A stronger difference in the clustering signal is found when considering galaxies of different colour (see, e.g., Li et al., 2006); at fixed M⋆ (or luminosity), red galaxies sensibly more clustered than blue galaxies. Galaxy colour is sensitive to the star formation rate (SFR) of model galaxies, which is regulated by a complex network of internal and external physical mechanisms. Therefore, clustering predictions for the active and passive galaxy populations represent a additional diagnostic to assess whether a given model reproduces the observational data in terms of star formation quenching as a function of the environment.
These results have been then confirmed and extended to higher redshift, thanks to redshift surveys such as the Galaxy And Mass Assembly (GAMA), VIMOS111Visible Multi Object Spectrograph Public Extragalactic Redshift Survey (VIPERS), the VIMOS-VLT Deep Survey (VVDS), the VIMOS Ultra Deep Survey (VUDS), the Deep Extragalactic Evolutionary Probe (DEEP2), the spectroscopic redshift Cosmic Evolution Survey (zCOSMOS) and showing that these conclusions generally hold up to z1 (see, among others, Coil et al., 2006; Pollo et al., 2006; Meneux et al., 2008, 2009; Marulli et al., 2013; Skibba et al., 2014; Farrow et al., 2015; Coil et al., 2017) and providing the first estimates of the 2pCF at z2 (Durkalec et al., 2018). Overall, there is still some debate about the redshift evolution of the amplitude of the 2pCF: results from DEEP2 and VVDS favour a small evolution of clustering amplitude with redshift, while VIPERS and zCOSMOS show negligible differences with respect to z0. It worth stressing the difficulties in comparing clustering measurements at different redshift, due to the different galaxy selection biases reflecting in the sampling of the different galaxy population. Moreover, the difference in clustering amplitudes of active and passive galaxies has been confirmed out to z2: Coil et al. (2017) studied the clustering as a function of SFR, and they found the same trends defined by colours at low-z, galaxies below the star-forming main sequence are more clustered than galaxies above it (see also Mostek et al., 2013).
In this paper, we exploit predictions from the latest version of the GAlaxy Evolution and Assembly (gaea), that couples an explicit partitioning of the cold gas into its molecular and neutral phases with an improved modelling of both cold gas accretion onto central Super-Massive Black Holes (SMBHs), and gas stripping in dense environments. We showed in our recent work (De Lucia et al., 2024; Xie et al., 2024) that this model realization correctly reproduces the observed fractions of quenched galaxies up to z3-4, while predicting number densities of massive quiescent galaxies at z3 that are the largest among recently published models. We will thus take advantage of the most recent determinations for the 2pCF at several redshifts to assess if our model is also able to reproduce both the global spatial distribution of galaxies in the LSS, and the distribution of galaxy populations split by their star formation activity. The latter test will represent a critical confirmation of the activity of gaea to correctly grasp the physical mechanisms responsible for the regulation of star formation in galaxies.
Most of our previous gaea realizations have been run on dark matter merger trees extracted from the Millennium Simulation (Springel et al., 2005, MS hereafter). This realization is based on a cosmological model whose parameters are offset with respect to more recent measurements – particularly for the normalisation of the density fluctuations at the present day () and . In this paper, we will introduce a novel gaea realization, run on the p-millennium simulation (Baugh et al., 2019, PMS hereafter), featuring a bigger cosmological volume, an updated cosmology, a better mass resolution and a finer time sampling. In particular, we will discuss the redshift evolution of key statistical quantities like the GSMFs or the AGN LFs as well as of galaxy clustering, as a function of stellar mass and star formation activity.
This paper is organized as follows. In Section 2 we will present the new gaea realization run on merger trees extracted from the PMS simulation. In Section 3 we compare these predictions with previous realizations of the same model run on different cosmological simulations. In Section 4 we discuss the predicted clustering of model galaxies at different redshift and as a function of different galaxy properties. Finally, we present and discuss our conclusions in Section 5.
2 N-body simulations and the galaxy formation model
Log() | |||||||||
MS | 0.75 | 0.25 | 0.0455 | 0.9 | 1.0 | 73 | 500 | ||
MSII | 0.75 | 0.25 | 0.0455 | 0.9 | 1.0 | 73 | 100 | ||
PMS | 0.693 | 0.307 | 0.04825 | 0.8288 | 0.9611 | 67.77 | 542.16 |
Parameter | Value | Meaning |
Star Formation efficiency | ||
0.13 | Reheating efficiency | |
0.10 | Ejection efficiency | |
0.68 | Reincorporation efficiency | |
0.6 | Hot gas Black Hole accretion efficiency | |
6.0 | Cold gas angular momentum loss efficiency | |
0.09 | Black hole accretion rate from reservoir | |
3.2 | Quasar wind efficiency | |
3.0 | fraction of ISM added to the BH reservoir as a consequence of AGN-driven outflows | |
rpstime | 400 | Timescale of hot gas ram-pressure stripping |
kesislow | 1.4 | Ratio between specific angular momentum of gas cooling through ‘slow mode’ and that of the halo |
kesirapid | 3.0 | Ratio between specific angular momentum of gas cooling through ‘rapid mode’ and that of the halo |
2.1 Numerical Simulations
In this work, we will focus on predictions from a gaea realization coupled with merger trees extracted from the p-millennium N-body simulation Baugh et al. (PMS 2019). Moreover, we will compare our results against predictions from two other gaea runs, based on the millennium Simulation (Springel et al., 2005, MS) and millennium ii Simulation (Boylan-Kolchin et al., 2009, MSII). These runs corresponds to the predictions at the basis of our recent work (De Lucia et al., 2024; Xie et al., 2024). Details on the assumed cosmologies, box size and particle resolution for each N-body simulation are listed in Table 1.
The PMS follows the evolution of the matter distribution in a volume of 8003 Mpc3, employing roughly 128 billion particles to trace the assembly of the Dark Matter Haloes (DMHs). In terms of mass resolution (), the PMS provides intermediate results to the MS and MSII. Moreover, it also provides an updated cosmology, based on the first year results from the Planck satellite (Planck Collaboration XVI, 2014), with respect to the MS and MSII, that were run assuming cosmological parameters consistent with first year results of the WMAP satellite (Spergel et al., 2003).
The initial conditions for the PMS have been generated at z=127 using second order Lagrangian perturbation theory (see Jenkins, 2013, for more details), and the simulation has been run using a reduced memory version of the N-body code gadget (Springel et al., 2005). Halo and Sub-Halo catalogues have been extracted on the fly using a classical FOF algorithm and the substructure finder code subfind (Springel et al., 2001): halos and sub-haloes information have been saved on a regular grid of 271 output times (i.e. snapshots), equally spaced in expansion factor. This has to be compared with the 64-68 snapshots of the MS and MSII over the same redshift range. The gain in temporal (redshift) resolution is particularly relevant at z4 where the PMS has 100 snapshots against the of the MS/MSII. In order to run GAEA on the PMS volume, we have constructed subhalo based merger trees using the same approach described in Springel et al. (2005). We refer to this paper for a detailed description of the algorithm. In summary, for any given subhalo, we identify a unique descendant by finding all subhalos at the subsequent output that contain its most bound particles. We then give higher weight to those particles that are more tightly bound to the subhalo under consideration, and select as descendant the subhalo containing a larger fraction of the most bound particles. In this way, we aim at tracing the fate of the inner regions of the substructure, that might survive for longer time after it has been accreted onto a larger structure. Following (Springel et al., 2005), we search for descendants in the two following outputs to deal with small subhalos that fluctuate below and above the assumed detection threshold (20 bound particles). Once a unique descendant has been identified for all subhalos in the simulated volume and for all available outputs, links to all progenitors are automatically established, and merger trees can be built. We adopt the same merger tree structure defined for the Millennium Simulation suite and shown in Fig. 11 of (Springel et al., 2005). The merger trees adopted in this work have been constructed using half of the available outputs. While reducing the number of redshifts available, this choice reduces the memory and computational requirements of our GAEA runs. In a forthcoming paper (Cantarella et al., in preparation), we will show that results are not significantly different from those obtained using merger trees constructed considering all available snapshots.
In Fig. 1 (left panel), we compare the DM halo mass functions (HMFs) derived from the subfind subhalo group catalogues for the MS (shaded areas), MSII (dashed lines) and PMS (point with poissonian errorbars), in the redshift range 0¡z¡5. The MS and MSII agree well over the mass range where both simulations have enough volume and resolution. On the other hand, the PMS shows less structures at high redshift and a faster evolution of the mass function, that is very close to that computed from the MS at z0. In order to understand the origin of these differences we consider in Fig. 1 (right panel) the evolution of the HMF shape for varying cosmological parameters. We evaluate analytic fits to the HMF at by assuming the Tinker et al. (2008) model and using the HMFcal tool developed by Murray et al. (2013). In all panels, the dashed black lines refer to HMFs predicted assuming the MS cosmology. The red solid lines in the left panels correspond to HMFs in the PMS cosmology. In the other panels, we show the HMFs predicted by varying only a few selected parameters with respect to the MS cosmology. In particular we consider MS models with (blue solid lines), with and (cyan dotted lines), with (green solid lines). This comparison highlights that the differences between MS and PMS are mainly driven by the lower value of used in PMS . The difference in and result in (smaller) opposite effects, enhancing and contrasting the trends induced by the different , respectively.
2.2 gaea
In this study we consider predictions from our latest rendition of the gaea semi-analytic model (SAM). The approach starts from a statistical description of the assembly of the Large Scale Structure (i.e. a merger tree) that can be either analytically derived or extracted from numerical experiments. Galaxies are assumed to form within DM haloes, under the effect of a complex network of physical processes responsible for energy and mass exchanges among different baryonic components. SAMs follow the evolution of galaxy populations by numerically solving a system of differential equations, aiming at describing specific physical mechanisms, using parametrizations derived from empirical, numerical or theoretical arguments. The parameters involved are calibrated against a well defined set of observational constraints, by sampling efficiently the parameter space. The main advantage of the SAM approach lies in its reduced computational demand, with respect to an hydro-dynamical simulations: they are an optimal tool for studying galaxy evolution over large cosmological volumes, with enough flexibility to extensively test different physical models, explore their associated parameter space, and therefore the impact of individual physical mechanisms to set the observed properties of different galaxy populations.
gaea builds from the model originally presented in De Lucia & Blaizot (2007), but now includes several improvements and further developments. In particular: (a) the modelling of chemical enrichment, based on a non-instantaneous treatment of the ejection of gas, metals and energy that accounts for the mass dependence of stellar lifetimes (De Lucia et al., 2014); (b) an improved treatment for stellar feedback, partially based on results of high-resolution hydrodynamical simulations (Hirschmann et al., 2016); (c) a detailed tracing of the evolution of angular momentum (Xie et al., 2017). The latest rendition of the model has been presented in De Lucia et al. (2024, DL24 hereafter) and merges independently developed versions, namely (d) the Xie et al. (2017) explicit partition of the cold gas into its atomic and molecular components; (e) the Xie et al. (2020) treatment for the non-instantaneous stripping of cold and hot gas in satellites galaxies; and (f) the Fontanot et al. (2020) modelling for cold gas accretion onto super-massive Black Holes (SMBH) and the onset of AGN-driven outflows. This new version of the gaea model has been calibrated on the evolution of the galaxy stellar mass function (GSMF), on the evolution of the AGN luminosity function (LFs) and on the local HI and H2 mass function. DL24 shows that it correctly reproduces the evolution of the fraction and number densities of quenched galaxies up to , providing the largest number densities of massive galaxies at among similar theoretical models, predicting the rise of the first massive quenched galaxies at (Xie et al., 2024). A more throughout analysis of the properties of the highest-z galaxies, especially in terms of their predicted UV luminosities as seen by JWST, will be presented in forthcoming work (Cantarella et al., in preparation). We stress that the latest gaea rendition preserves the good agreement with a variety of observational constraints that have been explored in recent years, such as the evolution of the mass-metallicity relations (Fontanot et al., 2021, both for the gaseous and stellar phases - see also Fig. 10), the GSMF and the cosmic star formation rate density up to the highest redshifts for which observations are available (Fontanot et al., 2017).
3 Convergence of model predictions on different simulations
In order to appreciate the effect of a change in the cosmological parameters and resolution on the predictions of our model, we consider gaea runs over merger trees extracted from the PMS , MS and MSII. All predictions we discuss in the following sections are based on model realization using the same values for the relevant physical parameters involved in the SAM definition. These parameters have been calibrated on the MS trees and correspond to the model presented in DL24. We list the most relevant in Table 2. In the following, we will show model predictions convolved with an estimated error on stellar and gaseous masses of 0.25 dex, thereby taking the Eddington bias (Jeffreys, 1938; Eddington, 1940) into account. The uncertainty considered represents a conservative choice especially at high-z. We contrast model predictions from the different simulations with compilations of observational data used in our previous works. In this section (Fig. 2), we focus on a basic set of physical properties, whose evolution is crucial for understanding the process of galaxy assembly, like the Galaxy Stellar Mass Function (GSMF - left panel) from z0 to z3 and the AGN bolometric Luminosity Function (LF, right panel) from z0 to z4. Those are among the main observables used to calibrate our model. In Appendix A, we include additional observational constraints considered for model calibration, plus several predictions we consider in our recent work.
It is remarkable that predictions corresponding to realizations run on different simulations show such a high level of consistency. This shows that our latest rendition of the gaea model exhibits an improved convergence with respect to previous versions. This is remarkable for the predicted evolution of the AGN-LF: the space density of faint AGNs is very sensible to the number of trigger events (i.e. mergers) and in the past this has limited the convergence of our model to relatively bright luminosities (see e.g. Fig. A2 in Fontanot et al. 2020). The latest gaea version shows a good convergence between the MS and the PMS runs down to L erg/s at z¡1, and down to L erg/s at higher redshifts; predictions from the PMS and MSII agree even better around the knee of the LF, while at brighter luminosities, the MSII lacks volume to correctly reproduce the rarer sources. For all other statistical estimators the level of convergence between the different runs is high and in most cases predictions based on different simulations are virtually indistinguishable. We interpret this effect as a result of the interplay between the improved modelling of star formation, based on the partition of the cold gas as in Xie et al. (2017), and the new prescription for cold gas accretion onto the central SMBHs (Fontanot et al., 2020). Both physical mechanisms act as regulators of the total amount of cold gas available for star formation, chemical enrichment and SMBH accretion. In fact, versions of the model implementing just one of the two physical mechanisms (and run on different simulation) do not achieve the same level of convergence over such a wide range of redshift and physical properties, as that presented here.
4 Clustering Analysis
4.1 Redshift evolution of the 2-point correlation function
We then consider the predicted correlation functions for our model universes at different redshifts. In the following, we will focus on the projected 2-point correlation function ; this choice makes the comparison with observational data straightforward, and does not have to rely on de-projection algorithms. We study the evolution of the projected correlation function using estimates from several surveys: going from low-z to higher redshifts we compare against results from SDSS (Li et al., 2006, z0.1), GAMA (Farrow et al., 2015, z¡0.2), VIPERS (Marulli et al., 2013, ), DEEP2 (Li et al., 2006, z0.9), VVDS (Meneux et al., 2008, z0.9), zCOSMOS (zCOSMOS Meneux et al., 2009, z0.9) and VUDS (Durkalec et al., 2018, z2.7). We collect all these observational measurements in Fig. 3 and we contrast them with the prediction from gaea realizations based on the MS (black dashed lines) or the PMS (red solid lines). We compute from gaea using the python corrfunc package (Sinha & Garrison, 2020); model galaxy samples have been extracted from the simulated volumes considering all galaxies in the stellar mass range indicated in the legend, i.e. we do not attempt to replicate selection criteria specific of the different surveys considered. In our analysis we include an estimate for the typical error on M⋆ by convolving the intrinsic predictions of the model with a log-normal error distribution of amplitude 0.25 dex. We check that our main conclusions do not depend on this choice: as predictions based on intrinsic M⋆ provide a consistent picture.
Overall, the level of agreement between theoretical and observed correlation functions is impressive over the entire redshift range considered. Models run on the MS and PMS agree well with each other, but for a tendency of the MS runs to predict systematically stronger correlations. The effect is more pronounced at small stellar masses and close separations, and we ascribe it to the larger value of used in the MS. At z (Fig. 3, upper left panel), theoretical are very close to SDSS constraints for , but they tend to underestimate the correlation signal measured for more massive galaxies. Similar conclusions hold at intermediate redshifts () for VIPERS (Fig. 3, upper right panel), but the underprediction in the highest mass bin reduces considerably as the redshift increases. The situation is slightly more complex at z0.9, where different surveys suggest slightly different normalization for . This could be due to the different surveyed volume or to the presence of overdensities (like in the case of the zCOSMOS sample, (Meneux et al., 2009)). We explore the impact of cosmic variance by considering 125 independent and disjointed subvolumes in the PMS , and computing their corresponding 2pCF. The yellow areas in Fig. 3 highlight the range covered by these independent realizations: smaller amplitudes are indeed possible, especially at the largest separations. Nonetheless, gaea predictions lie within the clustering measurements obtained from different surveys at z0.9 (Fig. 3, lower left panel) and the agreement with data is good to the smallest separations measured. The predicted evolution of the normalization at z1 is rather small: in most cases can be well approximated by a single power law over a wide separation scale; gaea also reproduces the observed flattening of the relation at low-z for the smallest separations probed by the SDSS. Moving to higher redshifts (z), a larger difference in the ) normalization can be appreciated with respect to the VUDS data (Fig. 3, lower right panel): uncertainties in the observational determination are typically larger than at lower redshift, and gaea predictions are still consistent, although systematically larger, than the data. Cosmic variance can also have a large impact here (the VUDS volume correspond to 1.75 107 Mpc3, to be compared with the 4 106 Mpc3 of the PMS subvolumes). We thus conclude that gaea is are able to correctly capture the observed evolution of the correlation strength up to z3.5: this represents a major success for our latest gaea model release.
4.2 Colour/SFR dependence of the 2-point correlation function
As we already discussed, several authors (see e.g. Wang et al. 2013) pointed out that the clustering strength depends not only on , but also on the star formation activity of galaxy population, with red and quiescent galaxies being more clustered than their blue and star forming counterparts with similar . In Fig. 4, we compare model predictions with SDSS data at z0.1 (Li et al., 2006) and data at 0.9 (Coil et al., 2017) obtained combining the PRIsm MUlti-object Survey (PRIMUS Coil et al., 2011) with DEEP2. At both redshifts we consider model predictions corresponding to quiescent and star-forming subsamples, using separation thresholds based on their specific SFR (sSFR = SFR/M⋆), for consistency with our previous work (De Lucia et al., 2024). At z0.1, model galaxies are classified using a sSFR= 10-11 yrs-1 separation. This is formally different than the choice for SDSS galaxies, which have been split according to their g-r colour: however, we checked that using a classifications based on galaxy colours (e.g. as in Springel et al. 2018) does not qualitatively change our main results: gaea realization reproduces well the stronger clustering of red galaxies with respect to blue sources, and in particular the relative clustering amplitudes between the two populations. Models over-predict the clustering strength for red galaxies at intermediate masses, irrespectively of the assumed cosmology. Overall, the clustering strength of blue galaxies is better reproduced over the entire mass range considered. At large masses, the model under-predict the observed clustering signal, but the relative clustering between the red and blue population is still well represented. At higher redshift, we contrast clustering estimates from the PRIMUS and DEEP2 surveys with gaea predictions by splitting model galaxies into a quiescent and star forming subsamples, using the same (redshift-dependent) sSFR threshold employed in the data (Coil et al., 2017):
The agreement between data and model prediction in the single mass bin available (Log(M]) ) is very good, both for the red and blue populations. Overall, the results presented in Fig. 4 indicate that the treatment for physical mechanism responsible for quenching in gaea correctly reproduces not only the overall assembly and evolution of the quenched galaxy population (De Lucia et al., 2024; Xie et al., 2024), but also their spatial distribution in the cosmic web. This confirms that the model correctly capture the relative role of internal and environmental processes in shaping the galaxy properties in different environments.
4.3 2-point correlation function for HI selected galaxies
In the previous sections, we study the 2pCF of galaxies selected on the basis of their stellar content. A complementary analysis involves the study of the cold gas content of galaxies, through the distribution of HI selected galaxies. Gas distribution in galaxies at low redshift can be mapped using the 21 cm HI hyper-fine emission line. In particular, the Arecibo Fast Legacy ALFA Survey (ALFALFA; Giovanelli et al. 2005) samples more than 5000 galaxies out to z0.05, enabling the study of their clustering properties Papastergis et al. (2013), showed that HI selected galaxies exhibit a lower amplitude with respect to stellar mass selected samples. Guo et al. (2017) used the 70% complete ALFALFA catalogue and showed that the 2pCF of of HI-selected galaxies depends strongly on the HI mass, i.e. galaxies with higher HI reservoirs are more strongly clustered on scales above a few Mpc than their counterparts with lower HI content. We compare Guo et al. (2017) estimates at different HI-mass (MHI) thresholds with gaea predictions in Fig. 5; we find a general good agreement in the clustering amplitude and, most importantly, in its relative evolution as a function of MHI. The larger discrepancies are seen at the high-MHI end, although data and predictions are consistent within the statistical errorbars. This result is similar to those based on M⋆ selected galaxies, although the reference samples refer to different model galaxy populations, as most massive galaxies at in gaea have low gas fractions.
5 Discussion and conclusions
In this paper, we present predictions of the latest version of the gaea model, run on merger trees extracted from the p-millennium simulation. We compare model results against a compilation of observational measurements of galaxy clustering as a function of stellar mass and star formation rate, covering a wide redshift range from the local Universe up to the highest redshift when this measurement has been possible (i.e. z3.5). Our model shows a satisfactory agreement with the observed projected 2pCF, of comparable or superior quality with respect to independent semi-analytic codes (L-galaxies Guo et al. 2011; Henriques et al. 2017) and hydro-dynamical simulations (Springel et al., 2018; Artale et al., 2017). At z0, gaea results are close to those predicted from the TNG hydro-simulation suite (Springel et al., 2018): a good agreement with data is found at intermediate-to-low stellar mass bins and at the smallest scales, while the 2pCF in the highest mass bin has a smaller amplitude with respect to observational estimates. Similar results have been also found for predictions from the EAGLE simulation (Artale et al., 2017). Predictions from the L-galaxies model are able to better reproduce the amplitude in the highest-mass bin, but they either overpredict (Guo et al., 2011) or underpredict (Henriques et al., 2017) the clustering of intermediate-to-low mass galaxies. Overall, our model does not predict a significant evolution of the clustering strength with redshift, in good agreement with independent theoretical predictions (Artale et al., 2017) and observational measurements up to z1 (Marulli et al., 2013). At higher redshift, data seem to suggest a more significant evolution of the clustering amplitude than model predictions, although these are still in agreement with data within 1- errorbars and cosmic variance in the observed samples might play a relevant role. More data at z2 are definitely needed to clarify if the discrepancy between available measurements and model predictions is significant.
We estimate the clustering length corresponding to our mock samples, by considering the value corresponding to a 2pCF . gaea predicts values in the range 4 to 6 h-1 Mpc, mildly increasing with stellar mass. These values are in good agreement with observational constraints for intermediate-to-low mass galaxies (see e.g. Foucaud et al. 2010; Wake et al. 2011; Lin et al. 2012), but underestimate the expected values at M1011 h-1 M⊙ (typically larger than 10 h-1 Mpc). This is a reflection of the underestimate of the clustering amplitude in the most massive stellar mass bin that we discussed in Sec. 4. At fixed stellar mass, we find only a weak evolution of with redshift, in agreement with the numerical results from Artale et al. (2017).
If we split our model galaxies in star-forming and quiescent samples, we find that gaea correctly reproduces the increased clustering of red quiescent galaxies with respect to blue star forming systems. At M1011 h-1 M⊙, the agreement with data for the blue population is excellent, while red galaxies show a clustering excess at intermediate stellar masses. This shows that the modelling of physical mechanisms responsible for quenching in gaea (both internal and environmental) is able to reproduce the observed clustering strength and its variation with stellar mass. This is not a trivial success for our model, and it shows that our approach is not only able to recover the global properties of galaxy populations in a statistical sense, but also their spatial distribution in the LSS. For the most massive galaxies (larger than 1011 h-1 M⊙), both the clustering amplitude of the quiescent and star-forming populations are underestimated. This is consistent with the underestimate we find for the total population, and represents the main tension we found.
Most of the model galaxies contributing to the most massive bin are central galaxies. The underprediction of clustering amplitude for these galaxies highlights a mismatch between the predicted and expected mass of the host DMHs: gaea galaxies tend to live in DMHs slightly less massive than what suggested by the clustering analysis. On the other hand, at low-stellar masses, that are dominated by the satellite population, the good agreement between data and predictions for the 1-halo term (i.e. at scales 2-3 Mpc) suggests that the satellite distribution within halos is correctly reproduced. This implies that our treatment for the evolution of satellite galaxies as in Xie et al. (2020) is effective in recovering the main properties of this population, both in terms of stellar mass and star formation activity. Conversely, the excess of clustering signal at low-masses for the 2-halo term ((i.e. at scales 2-3 Mpc) suggest that gaea is predicting a larger than expected number of satellites in massive haloes: this can be related to our treatment of the merging times for orphan galaxies (i.e. satellite galaxies whose substructure has fallen below the resolution limit of the simulation).
In summary, the latest gaea version shows a remarkable level of agreement for model predictions run on mergers trees extracted from simulation of different resolution and (slightly) different CDM cosmological parameters. This result is driven by the interplay between the new treatment for AGN (and its feedback on the host galaxies) and the explicit partitioning of cold gas into its atomic and molecular components. Both physical mechanisms act as regulators of the cold gas in model galaxies. gaea reproduces well the observed dependence of galaxy clustering on stellar mass and star formation activity. Moreover, model predictions are also consistent with the clustering measurements corresponding to HI-selected galaxies. This is an important confirmation that our treatment of the mechanisms acting on the baryonic component of galaxies, and satellites in particular, is able to capture the relevant physical dependencies. gaea predicts a small redshift evolution of the amplitude of the 2pCF: this is consistent with available data up to z3.5, which we interpret as another indication that our model is able to correctly connect galaxies with their environment. At higher redshifts, our model tends to under-predict the observational measurements. However, cosmic variance may play an important role at these redshifts and larger observational samples are needed to draw firmer conclusions in this redshift regime.
Acknowledgements.
An introduction to gaea , a list of our recent work, as well as datafile containing published model predictions, can be found at https://sites.google.com/inaf.it/gaea/home. We acknowledge the use of INAF-OATs computational resources within the framework of the CHIPP project (Taffoni et al., 2020) and the INAF PLEIADI program (http://www.pleiadi.inaf.it). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We thank Chen Li for sharing his observational estimates for the clustering signal in the SDSS and DEEP2 surveys. We acknowledge stimulating discussions with Andrea Biviano, Pierluigi Monaco and Emiliano Sefusatti. MH acknowledges funding from the Swiss National Science Foundation (SNSF) via a PRIMA grant PR00P2-193577 ‘From cosmic dawn to high noon: the role of BHs for young galaxies’.References
- Artale et al. (2017) Artale, M. C., Pedrosa, S. E., Trayford, J. W., et al. 2017, MNRAS, 470, 1771
- Baugh et al. (2019) Baugh, C. M., Gonzalez-Perez, V., Lagos, C. D. P., et al. 2019, MNRAS, 483, 4922
- Boylan-Kolchin et al. (2009) Boylan-Kolchin, M., Springel, V., White, S. D. M., Jenkins, A., & Lemson, G. 2009, MNRAS, 398, 1150
- Coil et al. (2011) Coil, A. L., Blanton, M. R., Burles, S. M., et al. 2011, ApJ, 741, 8
- Coil et al. (2017) Coil, A. L., Mendez, A. J., Eisenstein, D. J., & Moustakas, J. 2017, ApJ, 838, 87
- Coil et al. (2006) Coil, A. L., Newman, J. A., Cooper, M. C., et al. 2006, ApJ, 644, 671
- Cole & Kaiser (1989) Cole, S. & Kaiser, N. 1989, MNRAS, 237, 1127
- Cooray & Sheth (2002) Cooray, A. & Sheth, R. 2002, Phys. Rep, 372, 1
- Davis & Geller (1976) Davis, M. & Geller, M. J. 1976, ApJ, 208, 13
- De Lucia & Blaizot (2007) De Lucia, G. & Blaizot, J. 2007, MNRAS, 375, 2
- De Lucia et al. (2024) De Lucia, G., Fontanot, F., Xie, L., & Hirschmann, M. 2024, A&A, 687, A68
- De Lucia et al. (2014) De Lucia, G., Tornatore, L., Frenk, C. S., et al. 2014, MNRAS, 445, 970
- Durkalec et al. (2018) Durkalec, A., Le Fèvre, O., Pollo, A., et al. 2018, A&A, 612, A42
- Eddington (1940) Eddington, A. S., S. 1940, MNRAS, 100, 354
- Farrow et al. (2015) Farrow, D. J., Cole, S., Norberg, P., et al. 2015, MNRAS, 454, 2120
- Fletcher et al. (2021) Fletcher, T. J., Saintonge, A., Soares, P. S., & Pontzen, A. 2021, MNRAS, 501, 411
- Fontanot et al. (2021) Fontanot, F., Calabrò, A., Talia, M., et al. 2021, MNRAS, 504, 4481
- Fontanot et al. (2020) Fontanot, F., De Lucia, G., Hirschmann, M., et al. 2020, MNRAS, 496, 3943
- Fontanot et al. (2009) Fontanot, F., De Lucia, G., Monaco, P., Somerville, R. S., & Santini, P. 2009, MNRAS, 397, 1776
- Fontanot et al. (2017) Fontanot, F., Hirschmann, M., & De Lucia, G. 2017, ApJ, 842, L14
- Foucaud et al. (2010) Foucaud, S., Conselice, C. J., Hartley, W. G., et al. 2010, MNRAS, 406, 147
- Giovanelli et al. (2005) Giovanelli, R., Haynes, M. P., Kent, B. R., et al. 2005, AJ, 130, 2598
- Guo et al. (2017) Guo, H., Li, C., Zheng, Z., et al. 2017, ApJ, 846, 61
- Guo & White (2014) Guo, Q. & White, S. 2014, MNRAS, 437, 3228
- Guo et al. (2011) Guo, Q., White, S., Boylan-Kolchin, M., et al. 2011, MNRAS, 413, 101
- Haynes et al. (2011) Haynes, M. P., Giovanelli, R., Martin, A. M., et al. 2011, AJ, 142, 170
- Henriques et al. (2017) Henriques, B. M. B., White, S. D. M., Thomas, P. A., et al. 2017, MNRAS, 469, 2626
- Hirschmann et al. (2016) Hirschmann, M., De Lucia, G., & Fontanot, F. 2016, MNRAS, 461, 1760
- Jeffreys (1938) Jeffreys, H. 1938, MNRAS, 98, 190
- Jenkins (2013) Jenkins, A. 2013, MNRAS, 434, 2094
- Jenkins et al. (1998) Jenkins, A., Frenk, C. S., Pearce, F. R., et al. 1998, ApJ, 499, 20
- Jones et al. (2018) Jones, M. G., Haynes, M. P., Giovanelli, R., & Moorman, C. 2018, MNRAS, 477, 2
- Keres et al. (2003) Keres, D., Yun, M. S., & Young, J. S. 2003, ApJ, 582, 659
- Kormendy & Ho (2013) Kormendy, J. & Ho, L. C. 2013, ARA&A, 51, 511
- Li et al. (2006) Li, C., Kauffmann, G., Jing, Y. P., et al. 2006, MNRAS, 368, 21
- Li & White (2009) Li, C. & White, S. D. M. 2009, MNRAS, 398, 2177
- Lin et al. (2012) Lin, L., Dickinson, M., Jian, H.-Y., et al. 2012, ApJ, 756, 71
- Marulli et al. (2013) Marulli, F., Bolzonella, M., Branchini, E., et al. 2013, A&A, 557, A17
- McConnell & Ma (2013) McConnell, N. J. & Ma, C.-P. 2013, ApJ, 764, 184
- Meneux et al. (2009) Meneux, B., Guzzo, L., de la Torre, S., et al. 2009, A&A, 505, 463
- Meneux et al. (2008) Meneux, B., Guzzo, L., Garilli, B., et al. 2008, A&A, 478, 299
- Mostek et al. (2013) Mostek, N., Coil, A. L., Cooper, M., et al. 2013, ApJ, 767, 89
- Murray et al. (2013) Murray, S. G., Power, C., & Robotham, A. S. G. 2013, Astronomy and Computing, 3, 23
- Papastergis et al. (2013) Papastergis, E., Giovanelli, R., Haynes, M. P., Rodríguez-Puebla, A., & Jones, M. G. 2013, ApJ, 776, 43
- Planck Collaboration XVI (2014) Planck Collaboration XVI. 2014, A&A, 571, A16
- Pollo et al. (2006) Pollo, A., Guzzo, L., Le Fèvre, O., et al. 2006, A&A, 451, 409
- Sánchez et al. (2006) Sánchez, A. G., Baugh, C. M., Percival, W. J., et al. 2006, MNRAS, 366, 189
- Shen et al. (2020) Shen, X., Hopkins, P. F., Faucher-Giguère, C.-A., et al. 2020, MNRAS, 495, 3252
- Sinha & Garrison (2020) Sinha, M. & Garrison, L. H. 2020, MNRAS, 491, 3022
- Skibba et al. (2014) Skibba, R. A., Smith, M. S. M., Coil, A. L., et al. 2014, ApJ, 784, 128
- Spergel et al. (2003) Spergel, D. N., Verde, L., Peiris, H. V., et al. 2003, ApJS, 148, 175
- Springel et al. (2018) Springel, V., Pakmor, R., Pillepich, A., et al. 2018, MNRAS, 475, 676
- Springel et al. (2005) Springel, V., White, S. D. M., Jenkins, A., et al. 2005, Nature, 435, 629
- Springel et al. (2001) Springel, V., White, S. D. M., Tormen, G., & Kauffmann, G. 2001, MNRAS, 328, 726
- Taffoni et al. (2020) Taffoni, G., Becciani, U., Garilli, B., et al. 2020, arXiv e-prints, arXiv:2002.01283
- Tinker et al. (2008) Tinker, J., Kravtsov, A. V., Klypin, A., et al. 2008, ApJ, 688, 709
- Wake et al. (2011) Wake, D. A., Whitaker, K. E., Labbé, I., et al. 2011, ApJ, 728, 46
- Wang et al. (2013) Wang, L., Weinmann, S. M., De Lucia, G., & Yang, X. 2013, MNRAS, 433, 515
- Xie et al. (2024) Xie, L., De Lucia, G., Fontanot, F., et al. 2024, ApJ, 966, L2
- Xie et al. (2020) Xie, L., De Lucia, G., Hirschmann, M., & Fontanot, F. 2020, MNRAS, 498, 4327
- Xie et al. (2017) Xie, L., De Lucia, G., Hirschmann, M., Fontanot, F., & Zoldan, A. 2017, MNRAS, 469, 968
- Zehavi et al. (2002) Zehavi, I., Blanton, M. R., Frieman, J. A., et al. 2002, ApJ, 571, 172
- Zwaan et al. (2005) Zwaan, M. A., Meyer, M. J., Staveley-Smith, L., & Webster, R. L. 2005, MNRAS, 359, L30
Appendix A Additional model predictions
In this Appendix we show additional observational constraints to further explore the convergence of the latest gaea model, and to demonstrate that our previously published results hold in the latest rendition of our model. We consider in Fig. 6 the z0 distribution of sSFR in stellar mass bins and in Fig. 7 the evolution of the main sequence of star forming galaxies. In the latter figure, model star-forming galaxies have been selected as those with sSFR larger than 0.3/tH, where tH is the Hubble time at the redshift under consideration. Fig. 8 shows the z0 scaling relation between the mass of the central SMBH (MBH) and the mass of the spheroidal component (MBulge), while Fig. 9 shows the gas-phase and stellar mass-metallicity relations. In Fig. 10, we report the mass functions at for the neutral and molecular gas components. These distributions have been used for the calibration of the model on MS merger trees (Xie et al. 2017; De Lucia et al. 2024). Fig. 10 shows that for the HI and H2 mass function the consistency between the model realizations run on different simulations is higher (i.e. it extends to smaller gas masses) than previous versions of the model (see e.g. Fig. 7 in Xie et al. 2017).

### v2

@@
-Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model
-Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model
-We present results from the latest version of the GAlaxy Evolution and Assembly (gaea ) theoretical model of galaxy formation coupled with merger trees extracted from the P-Millennium Simulation (PMS ).
+Galaxy Assembly and Evolution in the P-Millennium simulation: galaxy clustering
+Galaxy Assembly and Evolution in the P-Millennium simulation: galaxy clustering
+We present results from the latest version of the GAlaxy Evolution and Assembly (gaea ) theoretical model of galaxy formation coupled with merger trees extracted from the Planck Millennium Simulation (PMS ).
 With respect to the Millennium Simulation, typically adopted in our previous work, the PMS provides a better mass resolution ( 10), a larger volume (8003 Mpc3) and assumes cosmological parameters consistent with latest results from the Planck mission.
-The model includes, at the same time, a treatment for the partition of cold gas into atomic and molecular (H2) components; a better treatment for environmental processes acting on satellite galaxies; an updated modelling of cold gas accretion on Super-Massive Black Hole, leading to the Active Galactic Nuclei (AGN) phenomenon and relative feedback on the host galaxy.
-We compare gaea predictions based on the PMS , with model realizations based on other simulations in the Millennium Suite at different resolution, showing that the new model provides a remarkable consistency in most statistical properties of galaxy populations.
-We interpret this as due to the interplay between AGN feedback and H2-based SFR (both acting as regulators of the cold gas content in model galaxies), as model version considering only one of the two mechanisms do not show the same level of consistency in the predictions.
+The model includes, at the same time, a treatment for the partition of cold gas into atomic and molecular (H2) components; a better treatment for environmental processes acting on satellite galaxies; an updated modelling of cold gas accretion on Super-Massive Black Holes, leading to the Active Galactic Nuclei (AGN) phenomenon and relative feedback on the host galaxy.
+We compare gaea predictions based on the PMS , with model realizations based on other simulations in the Millennium Suite at different resolution, showing that the new model provides a remarkable consistency for most statistical properties of galaxy populations.
+We interpret this as due to the interplay between AGN feedback and H2-based SFR (both acting as regulators of the cold gas content of model galaxies), as model versions considering only one of the two mechanisms do not show the same level of consistency.
 We then compare model predictions with available data for the galaxy 2-point correlation function (2pCF) in the redshift range 0¡z3.
-We show that gaea runs are able to correctly recover the main dependencies of the 2pCF as a function of stellar mass (M⋆), star formation activity, HI-content and redshift for M⋆ ¡ 1011 M⊙ galaxies.
+We show that gaea runs correctly recover the main dependencies of the 2pCF as a function of stellar mass (M⋆), star formation activity, HI-content and redshift for M⋆ ¿ 109 M⊙ galaxies.
 These results suggest that our model correctly captures both the distribution of galaxy populations in the Large Scale Structure and the interplay between the main physical processes regulating their baryonic content, both for central and satellite galaxies.
-At larger stellar masses gaea underpredicts the 2pCF amplitude, suggesting that model massive galaxies live in less massive dark matter haloes.
-The model predicts a rather small redshift evolution of the clustering amplitude up to z3, consistent with available observational evidence.
+The model predicts a small redshift evolution of the clustering amplitude, that results in an overprediction of z3 clustering strength with respect to the available estimates, but is still consistent with data within 1- uncertainties.
 Key Words.:
@@
 The 1-halo term is responsible for the 2pCF shape at 2 Mpc and it is statistically dominated by galaxy pairs belonging to the same halo (that is at least one of the two galaxies is a satellite).
-Overall, the largest contribution to the global 2pCF comes from galaxies around the knee of the GSMF, i.e.
-with stellar masses similar to the Milky-Way (Li & White, 2009).
-The small scale clustering is particularly sensitive to the number and distribution of satellite galaxies inside massive DMHs.
-Therefore, it provides important constraints to theoretical models of galaxy formation, as it critically depends on the treatment adopted for satellite galaxies and in particular their survival times.
-Therefore, it can be significantly affected by numerical resolution effects (Guo & White, 2014).
 Observationally, only the projected 2pCF - - is directly accessible, while the tridimensional 2pCF can be recovered via de-projection.
@@
 This realization is based on a cosmological model whose parameters are offset with respect to more recent measurements – particularly for the normalisation of the density fluctuations at the present day () and .
-In this paper, we will introduce a novel gaea realization, run on the p-millennium simulation (Baugh et al., 2019, PMS hereafter), featuring a bigger cosmological volume, an updated cosmology, a better mass resolution and a finer time sampling.
+In this paper, we will introduce a novel gaea realization, run on the planck millennium simulation (Baugh et al., 2019, PMS hereafter), featuring a bigger cosmological volume, an updated cosmology, a better mass resolution and a finer time sampling.
 In particular, we will discuss the redshift evolution of key statistical quantities like the GSMFs or the AGN LFs as well as of galaxy clustering, as a function of stellar mass and star formation activity.
@@
 Parameter | Value | Meaning |
-Star Formation efficiency | ||
+0.073 | Star Formation efficiency | |
 0.28 | Reheating efficiency | |
@@
 3.39 | fraction of ISM added to the BH reservoir as a consequence of AGN-driven outflows | |
-rpstime | 400 | Timescale of hot gas ram-pressure stripping |
-kesislow | 1.4 | Ratio between specific angular momentum of gas cooling through ‘slow mode’ and that of the halo |
-kesirapid | 3.0 | Ratio between specific angular momentum of gas cooling through ‘rapid mode’ and that of the halo |
+rpstime | 444 | Timescale of hot gas ram-pressure stripping |
+kesislow | 1.19 | Ratio between specific angular momentum of gas cooling through ‘slow mode’ and that of the halo |
+kesirapid | 1.38 | Ratio between specific angular momentum of gas cooling through ‘rapid mode’ and that of the halo |
 2.1 Numerical Simulations
-In this work, we will focus on predictions from a gaea realization coupled with merger trees extracted from the p-millennium N-body simulation Baugh et al.
-(PMS 2019).
+In this work, we will focus on predictions from a gaea realization coupled with merger trees extracted from the PMS N-body simulation (Baugh et al., 2019).
 Moreover, we will compare our results against predictions from two other gaea runs, based on the millennium Simulation (Springel et al., 2005, MS) and millennium ii Simulation (Boylan-Kolchin et al., 2009, MSII).
@@
 Once a unique descendant has been identified for all subhalos in the simulated volume and for all available outputs, links to all progenitors are automatically established, and merger trees can be built.
-We adopt the same merger tree structure defined for the Millennium Simulation suite and shown in Fig.
+We adopt the same merger tree structure defined for the MS suite and shown in Fig.
 11 of (Springel et al., 2005).
@@
 We stress that the latest gaea rendition preserves the good agreement with a variety of observational constraints that have been explored in recent years, such as the evolution of the mass-metallicity relations (Fontanot et al., 2021, both for the gaseous and stellar phases - see also Fig.
-10), the GSMF and the cosmic star formation rate density up to the highest redshifts for which observations are available (Fontanot et al., 2017).
+12), the GSMF and the cosmic star formation rate density up to the highest redshifts for which observations are available (Fontanot et al., 2017).
 3 Convergence of model predictions on different simulations
@@
 In the following, we will focus on the projected 2-point correlation function ; this choice makes the comparison with observational data straightforward, and does not have to rely on de-projection algorithms.
-We study the evolution of the projected correlation function using estimates from several surveys: going from low-z to higher redshifts we compare against results from SDSS (Li et al., 2006, z0.1), GAMA (Farrow et al., 2015, z¡0.2), VIPERS (Marulli et al., 2013, ), DEEP2 (Li et al., 2006, z0.9), VVDS (Meneux et al., 2008, z0.9), zCOSMOS (zCOSMOS Meneux et al., 2009, z0.9) and VUDS (Durkalec et al., 2018, z2.7).
+We study the evolution of the projected correlation function using estimates from several surveys: going from low-z to higher redshifts we compare against results from SDSS (Li et al., 2006, z0.1, maximum integration scale222The maximum integration scale represents the finite upper limit used for the integration of the real-space correlation function.
+=40 ), GAMA (Farrow et al., 2015, z¡0.2, =40 ), VIPERS (Marulli et al., 2013, , =30 ), DEEP2 (Li et al., 2006, z0.9, =40 ), VVDS (Meneux et al., 2008, z0.9, =20 ), zCOSMOS (zCOSMOS Meneux et al., 2009, z0.9, =20 ) and VUDS (Durkalec et al., 2018, z2.7, =20 ).
 We collect all these observational measurements in Fig.
-3 and we contrast them with the prediction from gaea realizations based on the MS (black dashed lines) or the PMS (red solid lines).
+3 and we contrast them with the prediction from gaea realizations based on the MS (black dashed lines), the MSII (dotted lines) or the PMS (red solid lines).
+Where not visible, the dotted lines (MSII) overlap with the dashed lines (MSI).
+We do not show predictions from the MSII for bins corresponding to M10 M⊙, as there are too few galaxies above this mass limit to provide a statistical estimate for the 2pCF.
 We compute from gaea using the python corrfunc package (Sinha & Garrison, 2020); model galaxy samples have been extracted from the simulated volumes considering all galaxies in the stellar mass range indicated in the legend, i.e.
 we do not attempt to replicate selection criteria specific of the different surveys considered.
-In our analysis we include an estimate for the typical error on M⋆ by convolving the intrinsic predictions of the model with a log-normal error distribution of amplitude 0.25 dex.
+In each plot (as well as in Fig.
+4 and 6), model estimates have been computed using the same as in the observed samples we compare with, but for the z0.9 panel.
+In this case we show model predictions corresponding to =40 (i.e.
+the value used in the DEEP2 survey).
+In our analysis we include an estimate for the typical uncertainty on M⋆ by convolving the intrinsic predictions of the model with a log-normal error distribution of amplitude 0.25 dex.
 We check that our main conclusions do not depend on this choice: as predictions based on intrinsic M⋆ provide a consistent picture.
-Overall, the level of agreement between theoretical and observed correlation functions is impressive over the entire redshift range considered.
-Models run on the MS and PMS agree well with each other, but for a tendency of the MS runs to predict systematically stronger correlations.
-The effect is more pronounced at small stellar masses and close separations, and we ascribe it to the larger value of used in the MS.
+We also explore the impact of uncertainties in the theoretical determination by computing the jackknife errors using the CosmoBolognaLib package (Marulli et al., 2016).
+We show these uncertainties for the PMS as the orange areas.
+We also assess the impact of cosmic variance on the comparison between data and model predictions: we randomly extract from the PMS 100 boxes with the same volume as the observational survey.
+The yellow areas in Fig.
+3 highlight the range covered by these sub-volumes.
+Overall, jackknife errors are relevant at large separations and for the most massive galaxies.
+Cosmic variance is relevant at the highest redshift considered and for small surveyed volumes, where it leads to errors that are slightly larger than those based on jackknife at intermediate scales.
+Overall, the level of agreement between theoretical and observed correlation functions is satisfactory over the entire redshift range considered.
+Models run on the MS, MSII and PMS agree well with each other, but for a slight tendency of the MS and MSII runs to predict systematically stronger correlations than PMS .
+The effect is more pronounced at small stellar masses and close separations, and we ascribe it to the larger value of used in the MS and MSII.
 At z (Fig.
-3, upper left panel), theoretical are very close to SDSS constraints for , but they tend to underestimate the correlation signal measured for more massive galaxies.
+3, upper left panel), theoretical are very close to SDSS constraints for .
+Above and below this mass range, model predictions slightly underestimate and over-estimate the measured clustering signal, respectively.
+The lower normalization of the 2pCF for galaxies more massive than 10 is connected with the overprediction of the high-mass end of the z0 GSMF (top left panel of Fig.
+2).
+In this mass range, our model galaxies lie above the Mvir versus M⋆ relation that can be inferred from data using a classical subhalo abundance matching technique (Behroozi et al., 2013; Moster et al., 2013), i.e.
+gaea massive galaxies tend to live in DMHs less massive than what can be inferred using halo occupation distribution statistics.
+For galaxies in the lowest mass bin considered, the overprediction of the clustering amplitude is mainly driven by the mass distribution of satellites and centrals in the smallest resolved DM haloes, an effect already discussed in Wang et al.
+(2013a).
+These authors showed that to reproduce at the same time the GSMF and 2pCF in the framework of an halo occupation distribution model, satellite galaxies should be less massive than centrals, on average, at fixed stellar mass.
+They also considered results from two independent semi-analytic models that struggled to reproduce these results - a problem that is shared also by our GAEA model.
+It is worth stressing that gaea lacks an explicit treatment for tidal stripping of stars from satellite galaxies, a process that could move model predictions in the right direction and whose impact we plan to study in future work.
+Moreover, at scales 2 Mpc orphan galaxies (i.e.
+satellite galaxies whose substructure has fallen below the resolution limit of the simulation) represent a relevant population and their orbital evolution and merging times estimates might be incorrect (see e.g.
+De Lucia et al.
+2010 for a discussion).
 Similar conclusions hold at intermediate redshifts () for VIPERS (Fig.
-3, upper right panel), but the underprediction in the highest mass bin reduces considerably as the redshift increases.
+3, upper right panel).
 The situation is slightly more complex at z0.9, where different surveys suggest slightly different normalization for .
 This could be due to the different surveyed volume or to the presence of overdensities (like in the case of the zCOSMOS sample, (Meneux et al., 2009)).
-We explore the impact of cosmic variance by considering 125 independent and disjointed subvolumes in the PMS , and computing their corresponding 2pCF.
-The yellow areas in Fig.
-3 highlight the range covered by these independent realizations: smaller amplitudes are indeed possible, especially at the largest separations.
 Nonetheless, gaea predictions lie within the clustering measurements obtained from different surveys at z0.9 (Fig.
-3, lower left panel) and the agreement with data is good to the smallest separations measured.
-The predicted evolution of the normalization at z1 is rather small: in most cases can be well approximated by a single power law over a wide separation scale; gaea also reproduces the observed flattening of the relation at low-z for the smallest separations probed by the SDSS.
-Moving to higher redshifts (z), a larger difference in the ) normalization can be appreciated with respect to the VUDS data (Fig.
-3, lower right panel): uncertainties in the observational determination are typically larger than at lower redshift, and gaea predictions are still consistent, although systematically larger, than the data.
-Cosmic variance can also have a large impact here (the VUDS volume correspond to 1.75 107 Mpc3, to be compared with the 4 106 Mpc3 of the PMS subvolumes).
-We thus conclude that gaea is are able to correctly capture the observed evolution of the correlation strength up to z3.5: this represents a major success for our latest gaea model release.
+3, lower left panel) and the agreement with data is acceptable to the smallest separations measured.
+The predicted evolution of the normalization at z1 is rather small: in most cases can be well approximated by a single power law over a wide separation scale; gaea also captures the observed flattening of the relation at low-z for the smallest separations probed by the SDSS.
+Moving to higher redshifts (z), a larger discrepancy of the ) normalization can be appreciated with respect to the VUDS data (Fig.
+3, lower right panel): gaea predictions are still consistent with observational measurements within the uncertainties, but tend to lie systematically above the data.
+Cosmic variance can also have a non-negligible impact here (the VUDS volume correspond to 1.75 107 Mpc3 roughly 1/175th of the PMS ).
+Overall, we can conclude that gaea is able to satisfactory reproduce the observed evolutionary trends of the correlation strength up to z3.5.
+This represents a non-trivial success for our latest gaea model release.
 4.2 Colour/SFR dependence of the 2-point correlation function
@@
 Wang et al.
-2013) pointed out that the clustering strength depends not only on , but also on the star formation activity of galaxy population, with red and quiescent galaxies being more clustered than their blue and star forming counterparts with similar .
+2013b) pointed out that the clustering strength depends not only on , but also on the star formation activity of galaxy population, with red and quiescent galaxies being more clustered than their blue and star forming counterparts with similar .
 In Fig.
-4, we compare model predictions with SDSS data at z0.1 (Li et al., 2006) and data at 0.9 (Coil et al., 2017) obtained combining the PRIsm MUlti-object Survey (PRIMUS Coil et al., 2011) with DEEP2.
-At both redshifts we consider model predictions corresponding to quiescent and star-forming subsamples, using separation thresholds based on their specific SFR (sSFR = SFR/M⋆), for consistency with our previous work (De Lucia et al., 2024).
+4, we compare model predictions with SDSS data at z0.1 (Li et al., 2006, =40 ) and data up to z1.0 (Coil et al., 2017, =40 ) obtained combining the PRIsm MUlti-object Survey (PRIMUS Coil et al., 2011) with DEEP2.
+At all redshifts, we consider model predictions corresponding to quiescent and star-forming subsamples, using separation thresholds based on their specific SFR (sSFR = SFR/M⋆), for consistency with our previous work (DL24).
 At z0.1, model galaxies are classified using a sSFR= 10-11 yrs-1 separation.
@@
 as in Springel et al.
-2018) does not qualitatively change our main results: gaea realization reproduces well the stronger clustering of red galaxies with respect to blue sources, and in particular the relative clustering amplitudes between the two populations.
-Models over-predict the clustering strength for red galaxies at intermediate masses, irrespectively of the assumed cosmology.
-Overall, the clustering strength of blue galaxies is better reproduced over the entire mass range considered.
-At large masses, the model under-predict the observed clustering signal, but the relative clustering between the red and blue population is still well represented.
+2018) does not qualitatively change our main results.
+Overall, gaea realizations show a stronger clustering signal of red galaxies with respect to blue ones, in qualitative agreement with observations.
+In more detail, our models reproduce well the clustering strength for red galaxies over the entire mass range considered.
+This is particularly true for the PMS realization, while some residual overestimate of the clustering signal holds at the lowest masses, and smallest scales.
+It is worth stressing that this is the range of scales and stellar masses where our model samples are dominated by orphan galaxies (that are typically associated with low levels of SFR and red colors).
+The clustering strength of blue galaxies is also reasonably well reproduced for intermediate masses (i.e.
+in the 109.5 ¡ M⋆/[] ¡ 1011 range).
+Outside this range, the clustering of star-forming galaxies shows the same trends discussed for the global population, thus suggesting that this galaxy population is driving the tensions between predictions and data reported in Sec.
+4.1 and Fig.
+3.
 At higher redshift, we contrast clustering estimates from the PRIMUS and DEEP2 surveys with gaea predictions by splitting model galaxies into a quiescent and star forming subsamples, using the same (redshift-dependent) sSFR threshold employed in the data (Coil et al., 2017):
-The agreement between data and model prediction in the single mass bin available (Log(M]) ) is very good, both for the red and blue populations.
-Overall, the results presented in Fig.
-4 indicate that the treatment for physical mechanism responsible for quenching in gaea correctly reproduces not only the overall assembly and evolution of the quenched galaxy population (De Lucia et al., 2024; Xie et al., 2024), but also their spatial distribution in the cosmic web.
-This confirms that the model correctly capture the relative role of internal and environmental processes in shaping the galaxy properties in different environments.
+The agreement between data and model predictions in the single mass bin available (Log(M]) ) is very good, both for the red and blue populations.
+The comparison presented in Fig.
+4 indicates that gaea realizations reproduce qualitatively the relative clustering amplitude between the red and blue populations.
+In order to better highlight this we consider in Fig.
+5 the ratio between the clustering amplitudes of the red and blue population at different redshifts for some of the available mass bins.
+Fig.
+5 shows that GAEA predictions reproduce qualitatively the observed trends, with deviations due to the discrepancies discussed in relation to Fig.
+4.
+The results presented in this section suggest that the treatment for the physical mechanisms responsible for quenching in gaea not only correctly reproduces the overall assembly and evolution of the quenched galaxy population (DL24, Xie et al.
+2024), but it is also broadly consistent with their expected distribution in the Large Scale Structure.
+Therefore, this analysis represents a further positive test suggesting that gaea correctly captures the relative role of internal and environmental processes in shaping the galaxy properties in different environments.
 4.3 2-point correlation function for HI selected galaxies
@@
 We compare Guo et al.
-(2017) estimates at different HI-mass (MHI) thresholds with gaea predictions in Fig.
-5; we find a general good agreement in the clustering amplitude and, most importantly, in its relative evolution as a function of MHI.
+(2017) estimates (=20 ) at different HI-mass (MHI) thresholds with gaea predictions in Fig.
+6; we find a general agreement in the clustering amplitude and, most importantly, in its relative evolution as a function of MHI.
 The larger discrepancies are seen at the high-MHI end, although data and predictions are consistent within the statistical errorbars.
@@
 5 Discussion and conclusions
-In this paper, we present predictions of the latest version of the gaea model, run on merger trees extracted from the p-millennium simulation.
+In this paper, we present predictions of the latest version of the gaea model, run on merger trees extracted from the PMS simulation.
 We compare model results against a compilation of observational measurements of galaxy clustering as a function of stellar mass and star formation rate, covering a wide redshift range from the local Universe up to the highest redshift when this measurement has been possible (i.e.
 z3.5).
-Our model shows a satisfactory agreement with the observed projected 2pCF, of comparable or superior quality with respect to independent semi-analytic codes (L-galaxies Guo et al.
+Our model shows a satisfactory agreement with the observed projected 2pCF, of comparable quality with respect to independent semi-analytic codes (L-galaxies Guo et al.
 2011; Henriques et al.
 2017) and hydro-dynamical simulations (Springel et al., 2018; Artale et al., 2017).
-At z0, gaea results are close to those predicted from the TNG hydro-simulation suite (Springel et al., 2018): a good agreement with data is found at intermediate-to-low stellar mass bins and at the smallest scales, while the 2pCF in the highest mass bin has a smaller amplitude with respect to observational estimates.
-Similar results have been also found for predictions from the EAGLE simulation (Artale et al., 2017).
-Predictions from the L-galaxies model are able to better reproduce the amplitude in the highest-mass bin, but they either overpredict (Guo et al., 2011) or underpredict (Henriques et al., 2017) the clustering of intermediate-to-low mass galaxies.
+At z0, gaea results are close to those predicted from the TNG hydro-simulation suite (Springel et al., 2018): a quite good agreement with data is found in the 10 ¡ M⋆ ¡ 10 stellar mass range and at the smallest scales, while the predicted 2pCT slightly under-predicts observational measurements for the most massive galaxies.
+Similar results have been also found for the EAGLE simulation (Artale et al., 2017).
+Predictions from the L-galaxies model are closer to the observed amplitude in the highest-mass bin, but they either overpredict (Guo et al., 2011) or underpredict (Henriques et al., 2017) the clustering of intermediate-to-low mass galaxies (i.e.
+10 M 10) At the lowest mass bin, our model predictions over-predict the clustering signal, similarly to what found by Guo et al.
+(2011).
 Overall, our model does not predict a significant evolution of the clustering strength with redshift, in good agreement with independent theoretical predictions (Artale et al., 2017) and observational measurements up to z1 (Marulli et al., 2013).
-At higher redshift, data seem to suggest a more significant evolution of the clustering amplitude than model predictions, although these are still in agreement with data within 1- errorbars and cosmic variance in the observed samples might play a relevant role.
-More data at z2 are definitely needed to clarify if the discrepancy between available measurements and model predictions is significant.
-We estimate the clustering length corresponding to our mock samples, by considering the value corresponding to a 2pCF .
-gaea predicts values in the range 4 to 6 h-1 Mpc, mildly increasing with stellar mass.
-These values are in good agreement with observational constraints for intermediate-to-low mass galaxies (see e.g.
+At higher redshift, data seem to suggest a more significant evolution of the clustering amplitude than model predictions, resulting in an overprediction with respect to the observed normalization of the 2pCF.
+Considering the observational uncertainties, model predictions are still consistent with observational data, and more data at z2 are needed to clarify if the highlighted discrepancy is significant.
+Fig.
+7 shows the dependence of the clustering lenght on galaxy stellar mass and its redshift evolution as predicted by our runs, and compares it to observational measurements.
+We compute from the intrinsic three-dimensional 2pCF (), directly computed from our model outputs, defined as the scale corresponding to .
+gaea predicts a significant increase of as a function of stellar mass.
+For M10 the predicted values are in reasonable agreement with observational constraints (see e.g.
 Foucaud et al.
@@
 2011; Lin et al.
-2012), but underestimate the expected values at M1011 h-1 M⊙ (typically larger than 10 h-1 Mpc).
-This is a reflection of the underestimate of the clustering amplitude in the most massive stellar mass bin that we discussed in Sec.
-4.
-At fixed stellar mass, we find only a weak evolution of with redshift, in agreement with the numerical results from Artale et al.
+2012; Marulli et al.
+2013), with a clear excess of the predicted value of r0 for galaxy stellar masses below 10.
+At fixed stellar mass, we find only a weak evolution of at , in agreement with the numerical results from Artale et al.
+(2017) and Springel et al.
 (2018).
+A significant discrepancy between model predictions and data is found at z2, with gaea predicting a factor of 2 increase of at fixed stellar mass, and data exhibiting a slower trend as a function of galaxy stellar mass evolution.
 If we split our model galaxies in star-forming and quiescent samples, we find that gaea correctly reproduces the increased clustering of red quiescent galaxies with respect to blue star forming systems.
-At M1011 h-1 M⊙, the agreement with data for the blue population is excellent, while red galaxies show a clustering excess at intermediate stellar masses.
-This shows that the modelling of physical mechanisms responsible for quenching in gaea (both internal and environmental) is able to reproduce the observed clustering strength and its variation with stellar mass.
-This is not a trivial success for our model, and it shows that our approach is not only able to recover the global properties of galaxy populations in a statistical sense, but also their spatial distribution in the LSS.
-For the most massive galaxies (larger than 1011 h-1 M⊙), both the clustering amplitude of the quiescent and star-forming populations are underestimated.
-This is consistent with the underestimate we find for the total population, and represents the main tension we found.
-Most of the model galaxies contributing to the most massive bin are central galaxies.
-The underprediction of clustering amplitude for these galaxies highlights a mismatch between the predicted and expected mass of the host DMHs: gaea galaxies tend to live in DMHs slightly less massive than what suggested by the clustering analysis.
-On the other hand, at low-stellar masses, that are dominated by the satellite population, the good agreement between data and predictions for the 1-halo term (i.e.
-at scales 2-3 Mpc) suggests that the satellite distribution within halos is correctly reproduced.
+Overall, the red population shows a better agreement with data over the entire galaxy mass range considered, while blue galaxies exhibit both an excess of the clustering signal for low stellar masses (M109 M⊙) and an underprediction of the 2pCF at large masses (M1011 M⊙).
+Both discrepancies are consistent with the tensions we find between model prediction and data for the global population.
+In general, gaea results align well with similar comparisons shown in Guo et al.
+(2011) and Springel et al.
+(2018) for an independent semi-analytic and hydro-dynamical model of galaxy formation.
+These models tend to better reproduce the properties of the star-forming population, while overpredicting the clustering of the red population at intermediate stellar mass (in particular, for M 109-1010 M⊙).
+Finally, we also compare our model predictions with clustering measurement for HI-selected galaxies in the ALFA survey and we find a good agreement over the entire mass range considered.
+Overall, the results discussed in this paper demonstrate that the proposed modelling of physical mechanisms responsible for gas consumption, recycling and quenching in gaea (both internal and environmental) are able to capture the main trends found for the observed clustering strength and its variation with stellar mass, SF activity and HI content.
+We stress that our model has not been calibrated to reproduce galaxy clustering, that thus represents a genuine model prediction: this is not a trivial success, and it shows that our approach is not only able to recover the global properties of galaxy populations in a statistical sense, but also their spatial distribution in the LSS.
+The general good agreement between model predictions and data for the 1-halo term (i.e.
+at scales 2-3 Mpc) for stellar masses down to M 109 M⊙, suggests that the satellite distribution within halos is correctly reproduced, at least down to this stellar mass limit.
 This implies that our treatment for the evolution of satellite galaxies as in Xie et al.
-(2020) is effective in recovering the main properties of this population, both in terms of stellar mass and star formation activity.
-Conversely, the excess of clustering signal at low-masses for the 2-halo term ((i.e.
-at scales 2-3 Mpc) suggest that gaea is predicting a larger than expected number of satellites in massive haloes: this can be related to our treatment of the merging times for orphan galaxies (i.e.
-satellite galaxies whose substructure has fallen below the resolution limit of the simulation).
-In summary, the latest gaea version shows a remarkable level of agreement for model predictions run on mergers trees extracted from simulation of different resolution and (slightly) different CDM cosmological parameters.
+(2020) is generally effective in recovering the main properties of this population, both in terms of stellar mass and star formation activity.
+The excess of clustering signal found for lower stellar masses can be understood as a combination of different effects.
+On the one hand, satellite galaxies residing in low-mass haloes are over-massive with respect to what is found in subhalo abundance matching models constrained to reproduce both the 2pCF and the GSMF (Wang et al., 2013a).
+The clustering signal excess for the 2-halo term ((i.e.
+at scales 2-3 Mpc) suggests that gaea is predicting a larger than expected number of satellites in massive haloes and this can be related to our treatment of the merging times for orphan galaxies.
+In summary, the latest gaea version shows a remarkable level of agreement for model predictions run on mergers trees extracted from simulations of different resolution and (slightly) different CDM cosmological parameters.
 This result is driven by the interplay between the new treatment for AGN (and its feedback on the host galaxies) and the explicit partitioning of cold gas into its atomic and molecular components.
 Both physical mechanisms act as regulators of the cold gas in model galaxies.
-gaea reproduces well the observed dependence of galaxy clustering on stellar mass and star formation activity.
-Moreover, model predictions are also consistent with the clustering measurements corresponding to HI-selected galaxies.
+gaea reasonably reproduces the observed dependence of galaxy clustering on stellar mass and star formation activity.
+Moreover, model predictions are also broadly consistent with the clustering measurements corresponding to HI-selected galaxies.
 This is an important confirmation that our treatment of the mechanisms acting on the baryonic component of galaxies, and satellites in particular, is able to capture the relevant physical dependencies.
@@
 At higher redshifts, our model tends to under-predict the observational measurements.
-However, cosmic variance may play an important role at these redshifts and larger observational samples are needed to draw firmer conclusions in this redshift regime.
+However, larger observational samples are needed to draw firmer conclusions in this redshift regime.
 Acknowledgements.
@@
 DiRAC is part of the National e-Infrastructure.
+We thank the anonymous referee for a careful reading of the manuscript that helped us improve the quality of the presentation of our results.
 We thank Chen Li for sharing his observational estimates for the clustering signal in the SDSS and DEEP2 surveys.
-We acknowledge stimulating discussions with Andrea Biviano, Pierluigi Monaco and Emiliano Sefusatti.
+We acknowledge stimulating discussions with Andrea Biviano, Alessandra Fumagalli, Federico Marulli, Pierluigi Monaco and Emiliano Sefusatti.
 MH acknowledges funding from the Swiss National Science Foundation (SNSF) via a PRIMA grant PR00P2-193577 ‘From cosmic dawn to high noon: the role of BHs for young galaxies’.References
@@
 - Baugh et al. (2019) Baugh, C. M., Gonzalez-Perez, V., Lagos, C. D. P., et al. 2019, MNRAS, 483, 4922
+- Behroozi et al. (2013) Behroozi, P. S., Wechsler, R. H., & Conroy, C. 2013, ApJ, 770, 57
 - Boylan-Kolchin et al. (2009) Boylan-Kolchin, M., Springel, V., White, S. D. M., Jenkins, A., & Lemson, G. 2009, MNRAS, 398, 1150
@@
 - De Lucia & Blaizot (2007) De Lucia, G. & Blaizot, J. 2007, MNRAS, 375, 2
+- De Lucia et al. (2010) De Lucia, G., Boylan-Kolchin, M., Benson, A. J., Fontanot, F., & Monaco, P. 2010, MNRAS, 406, 1533
 - De Lucia et al. (2024) De Lucia, G., Fontanot, F., Xie, L., & Hirschmann, M. 2024, A&A, 687, A68
@@
 - Guo et al. (2017) Guo, H., Li, C., Zheng, Z., et al. 2017, ApJ, 846, 61
-- Guo & White (2014) Guo, Q. & White, S. 2014, MNRAS, 437, 3228
 - Guo et al. (2011) Guo, Q., White, S., Boylan-Kolchin, M., et al. 2011, MNRAS, 413, 101
@@
 - Li et al. (2006) Li, C., Kauffmann, G., Jing, Y. P., et al. 2006, MNRAS, 368, 21
-- Li & White (2009) Li, C. & White, S. D. M. 2009, MNRAS, 398, 2177
 - Lin et al. (2012) Lin, L., Dickinson, M., Jian, H.-Y., et al. 2012, ApJ, 756, 71
 - Marulli et al. (2013) Marulli, F., Bolzonella, M., Branchini, E., et al. 2013, A&A, 557, A17
+- Marulli et al. (2016) Marulli, F., Veropalumbo, A., & Moresco, M. 2016, Astronomy and Computing, 14, 35
 - McConnell & Ma (2013) McConnell, N. J. & Ma, C.-P. 2013, ApJ, 764, 184
@@
 - Mostek et al. (2013) Mostek, N., Coil, A. L., Cooper, M., et al. 2013, ApJ, 767, 89
+- Moster et al. (2013) Moster, B. P., Naab, T., & White, S. D. M. 2013, MNRAS, 428, 3121
 - Murray et al. (2013) Murray, S. G., Power, C., & Robotham, A. S. G. 2013, Astronomy and Computing, 3, 23
@@
 - Wake et al. (2011) Wake, D. A., Whitaker, K. E., Labbé, I., et al. 2011, ApJ, 728, 46
-- Wang et al. (2013) Wang, L., Weinmann, S. M., De Lucia, G., & Yang, X. 2013, MNRAS, 433, 515
+- Wang et al. (2013a) Wang, L., De Lucia, G., & Weinmann, S. M. 2013a, MNRAS, 431, 600
+- Wang et al. (2013b) Wang, L., Weinmann, S. M., De Lucia, G., & Yang, X. 2013b, MNRAS, 433, 515
 - Xie et al. (2024) Xie, L., De Lucia, G., Fontanot, F., et al. 2024, ApJ, 966, L2
@@
 In Fig.
-10, we report the mass functions at for the neutral and molecular gas components.
+12, we report the mass functions at for the neutral and molecular gas components.
 These distributions have been used for the calibration of the model on MS merger trees (Xie et al.